English formulae pronunciation

No. Symbols  Reading
1 a_{1}  AY-sub-one
2 a_{j}  AY-sub-JEY
3 x^{⋆}  EX-star
4 x̄  EX-bar
5 x̂  EX-hat
6 x̃  EX-tilde
7 x_{ls}  EX-sub-EL-ES
8 x^{2}  EX-squared
9 x^{3}  EX-cubed
10 x^{n}  EX-to-the-EN-th-power
11 x^{T}  EX-transposed, transpose-of-EX
12 e^{x}  EE-to-power-EX, EE-to-the-EX-th-power
13 n!  EN-factorial
14 exp{x}  Exponential-function-of-EX
15 f(x)  EF-of-EX, EF-at-EX
16 f(a, x)  EF-of-AY-and-EX
17 δ[ε, x_{0}, f(x)]  DELTA-of-EPSILON-EX-sub-zero-and-EF-of-EX
18 f(x | z)  EF-of-EX-restricted-to-ZET
19 P_{⋆}^{−1}  Inverse-of-PEE-star
20 (\sqrt{2π})^{m}  Square-root-of-the-two-PAI-to-the-EM-th-power
21 αx  ALFA-times-EX

22

x/(x + b) or \frac{x}{x+b}

EX-divided-by-EX-plus-BEE, EX-divided-by-the-sum-of-EX-and-BEE

23

|P |^{1/2}

The-determinant-of-PEE-to-the-one-half-power, the-square-root of-the-determinant-of-PEE

24 E {x | x̄}  Conditional-mean-of-EX-restricted-to-the-EX-bar
25 A^{+}  Pseudo-inverse-of-AY

continued on next page

continued from previous page

No. Symbols  Reading

26

S_{n} = d_{1} + ⋯ + d_{n}

ES-sub-EN-is-the-sum-of-DEE-sub-one-through-DEE-sub-EN

27 \sqrt[n]{x}  The-EN-th-root-of-EX

28

P_{n} = μ_{1} ⋯ μ_{n}

PEE-sub-EN-is-the-product-of-MU-sub-one-through-MU-sub-EN

29

x ∈ A

EX-is-a-member-of-AY, EX-belongs-to-AY, EX-is-in-AY, EX-is-an-element-of-AY

30

A ⊂ B

AY-is-contained-in-BEE, BEE-contains-AY, AYis-a-subset-of-BEE

31 {x : P (x)}  The-set-of-EX-for-which-PEE-of-EX-is-true

32

{x ∈ X : P (x)}

The-set-of-EX-(to-come)-from-(the-set)-EX-uppercase-for-which-PEE-of-EX-is-true

33 A → B  AY-implies-BEE
34 ∀x ∈ C : P (x)  For-all-EX-in-SEE-PEE-of-EX-(is-true)
35 ∃x ∈ C : P (x)  For-at-least-one-EX-in-SEE-PEE-of-EX-(is-true)
36 1/2  One-half, one-second
37 1/3  One-third
38 5/60  Five-sixtieth
39 ¬A  Not-AY

40

∑_{j=1}^{m} f(j)

The-sum-of-EF-of-JEY-from-JEY-equals-one-to- (JEY-equals)-EM

41

∏_{j=1}^{m} f(j)

The-product-of-EF-of-JEY-from-JEY-equals-one-to-JEY-equals-EM

42

z = Ax + v

ZET-is-AY-times-EX-plus-VEE, ZET-equals-AY-times-EX-plus-VEE, ZET-is-the-equal-of-AY-times-EX-plus-VEE

43 x := z  Let-EX-be-the-equal-of-ZET

continued on next page

No. Symbols  Reading

44 2 + 2 = 4 Two and two are four. Two plus two equals four. Two and two make four. If you take 2 and 2, you get 4. (We add components (or summunds) and obtain their sum.)

45 4 − 2 = 2 Two from four are two. Four minus two equals two. If you take (subtract) 2 from 4, you get 2. (We subtract two from four and obtain the difference. 4 is the minuend; 2 is the subtrahend.)

46 2 × 2 = 4 Two times two is four. Twice two is four. Two times two equals four. Two muliplied by two is four. (We multiply factors and obtain the product.)

47 5 × 8 = 40  Five eights are fourty.

48 8 : 2 = 4 Two into eight is four. Two into eight goes four times. Eight divided by two equals (makes) four. (We divide eight by two and obtain the quotient. Two is the divisor. Eight is the dividend.)